# adF: a Novel System for Measuring Web Fingerprinting through Ads

### v2

@@ 6. Vulnerability and Accuracy metrics
6.1. Vulnerability Metrics
6.1.1. Theoretical Vulnerability ($\mathcal{TV}$)
@@
6.1.2. Measurable Vulnerability ($\mathcal{MV}$)
@@
6.2. adF system’s accuracy
Our adF system is essentially tackles a classification task, aiming to extract device fingerprints and determine their uniqueness. To assess its performance, we use a standard classification metric, accuracy ($\mathcal{A}$). Accuracy, in our scenario, represents the proportion of fingerprints correctly classified as unique (true positives) or non-unique (true negatives) by our system, relative to the total number of fingerprints ($N_{f}$). A high accuracy suggests that even without access to advertising ID information, our system can effectively assess the vulnerability of different device configurations to web fingerprinting.
@@
7. Methodology Limitations
Our dataset is limited in device coverage, so our claim of fingerprint uniqueness is confined to our dataset’s scope, and we cannot claim the global uniqueness of such fingerprints. This limitation stems primarily from our restricted budget for ad campaigns, amounting to 410,27 €. Also, ethical considerations led us to exclude almost half of the samples without an advertising ID in our dataset. Nevertheless, despite these constraints, our dataset’s size (see Section 8) is comparable to previous works in the literature (gomezBoix2018hiding, ; andriamilanto2021largescale, ). It confirms that adF provides high accuracy, establishing it as a reliable system for evaluating device configuration vulnerability to web fingerprinting. Furthermore, our results highlight the significant variability in protection among different device configurations.
We also acknowledge that our methodology faces different issues that impede collecting the required data from every single ad executing our Data Collector’s adTag. Some of these issues include: browsers running ad-blockers (adblock_plus, ) or disabling JavaScript (disable_javascript, ), outdated browsers, limited browser API support, network problems in the connection between the ad and our server, etc. Moreover, the runtime required by our adTag represents another source of losses. Our adTag runs as long as the webpage (mobile app) where the ad is rendered is open on the user’s device. Hence, once the user closes the webpage (mobile app), the adTag execution is stopped. Callejo et al. (callejo2017opportunities, ) estimated that 75% of ads displayed on PCs and mobile devices remain active for at least 15 and 8 seconds, respectively. Although our adTag has an overall estimated runtime in the order of hundreds of milliseconds (depending on the browser and OS), we must also consider the rendering time of the website itself. All the mentioned causes lead to an average loss of  19% ad impressions from web-pages and of  33% ad impressions from mobile apps by comparing the campaign’s number of impressions reported by the DSP and the total samples collected in our back-end server (See Table 1).
8. Results
@@ 8.1. Performance of our adF system
8.2. Vulnerability of browser-based device configurations to web fingerprinting
The values of $\mathcal{TV}$ and $\mathcal{MV}$ presented in Figure 2 serve to assess the vulnerability of each of the browser-based configurations to web fingerprinting. Given the similarity between $\mathcal{TV}$ and $\mathcal{MV}$, we will discuss results based solely on $\mathcal{MV}$ for the remainder of this section.
The first conclusion we extract from the results in Figure 2 is that mobile devices exhibit lower vulnerability to web fingerprinting when compared to desktop devices for a given browser (Chrome or Safari). For instance, in desktop devices running macOS, Safari’s $\mathcal{MV}$ is 5.5%, while in mobile devices running iOS, it is 3.4%. Likewise, Chrome offers lower $\mathcal{MV}$ in mobile devices compared to desktop devices, regardless of the OS.
@@ 8.2. Vulnerability of browser-based device configurations to web fingerprinting
Finally, when examining browsers individually, we find that Chrome ranks as the least secure option (except in the case of iOS). In contrast, Safari and Firefox provide notably superior protection against web fingerprinting. Note that we gather a few hundred samples from Firefox on both Windows and macOS, suggesting Firefox’s strong performance also on these operating systems. To better understand the outstanding performance of Chrome on iOS, we refer the reader to our analysis in Section 10. This result is due to the low cardinality and entropy of attributes accessible in Chrome for iOS devices, which does not occur in other OSes.
@@
[…]
8.3. Mobile apps
@@
Second, Android apps are more exposed to web fingerprinting compared to iOS. In particular, when considering $\mathcal{MV}$, we find that 40.7% of Android users and 22% of iOS users can be uniquely fingerprinted using mobile apps’ ads. If we extrapolate these results to the actual market share of Android and iOS (71,01% and 28,30%, respectively), we conclude that 35% of mobile devices can be uniquely identified using web fingerprinting techniques in mobile apps.
@@
9. Causes of Web Fingerprinting Vulnerability
@@
9.1. Reported attribute count ($|N|$)
@@
Longer vectors generally result in a higher probability of acquiring unique fingerprints. In other words, device configurations that provide a greater number of attributes are typically more vulnerable to web fingerprinting. However, this is not the sole factor influencing the resilience to web fingerprinting. For instance, in the case of desktop configurations, if we use Safari’s anti-fingerprint settings in Chrome samples on Windows (i.e., we do not consider in Chrome those attributes not reported by Safari), the $\mathcal{MV}$ ($\mathcal{TV}$) for {Chrome, Windows} devices become 61,46% (67,16%). This represents an important improvement compared to the original vulnerability measured for {Chrome, Windows} configuration, 95,74% (95.81%). However, it is still far from {Safari, macOS} results ($\mathcal{MV}$ = 5,48%, $\mathcal{TV}$ = 8.36%). This example shows that two device configurations reporting exactly the same attributes present significantly different web fingerprinting vulnerabilities. Hence, there must be other factors influencing the web fingerprinting vulnerability.
Table 2. Discrimination power of attributes. For each attribute and device configuration, we report: (i) whether the attribute is reported or not, (ii) the cardinality ($|S|$), (iii) the normalized entropy ($h$). We highlight in bold those attributes whose reporting we propose to block in our solution ($|S|>$ 25 and a $h\geq$ 0,10).
@@
UserAgent 6946 0,39 165 0,44 602 0,57 131 0,38 374 0,28 56 0,34 141 0,25 72 0,36 216 0,21 82 0,46 5050 0,73 106 0,33
CPU cores 8 0,02 2 0,01 4 0,03 4 0,05 24 0,14 9 0,22 13 0,14 4 0,12 24 0,13 9 0,23 6 0,02 2 0,03
Device memory 6 0,09 4 0,10 1 0,00 1 0,00 6 0,08 5 0,09 4 0,03 1 0,00 6 0,11 1 0,00 5 0,10 1 0,00
Screen: color depth 1 0,00 1 0,00 2 0,02 3 0,02 4 0,00 3 0,01 3 0,08 2 0,08 4 0,00 2 0,02
Screen: pixel left 4 0,00 1 0,00 1 0,00 2 0,00 1615 0,16 74 0,17 735 0,21 286 0,15 609 0,17 101 0,32
Screen: orientation angle 4 0,05 3 0,05 4 0,08 4 0,10 4 0,00 4 0,01 2 0,00 5 0,14 4 0,05 1 0,00 4 0,00 5 0,06
Screen: orientation type 4 0,05 3 0,05 4 0,08 4 0,10 4 0,00 4 0,00 3 0,00 5 0,14 4 0,07 1 0,00
Battery status: charging 3 0,04 3 0,05 1 0,00 1 0,00 3 0,03 3 0,05 3 0,08 1 0,00 3 0,07 1 0,00
Simultaneous touch points 4 0,02 1 0,00 3 0,00 4 0,00 19 0,02 6 0,05 2 0,00 3 0,08 8 0,09 2 0,00 2 0,03 2 0,01
Media devices 39 0,17 6 0,01 5 0,08 17 0,10 66 0,14 8 0,16 8 0,05 12 0,10 43 0,21 28 0,21 22 0,18 5 0,08
Languages 2748 0,31 22 0,13 27 0,15 39 0,18 1638 0,20 55 0,24 822 0,36 43 0,18 1355 0,39 40 0,25 285 0,17 42 0,20
PDF viewer enabled 3 0,01 2 0,05 3 0,08 3 0,08 3 0,04 3 0,01 3 0,01 3 0,09 3 0,08 3 0,09
User Permissions state 196 0,13 1 0,00 6 0,18 22 0,20 113 0,13 21 0,17 58 0,15 19 0,19 155 0,22 16 0,25
available height 456 0,32 41 0,31 30 0,29 32 0,25 549 0,26 79 0,42 582 0,52 310 0,45 590 0,44 206 0,63 227 0,33 20 0,24
available left 1 0,00 1 0,00 1 0,00 3 0,00 221 0,04 17 0,04 199 0,07 30 0,01 166 0,14 60 0,27
available top 1 0,00 1 0,00 1 0,00 3 0,00 484 0,03 13 0,02 181 0,18 13 0,11 170 0,12 38 0,28
available width 445 0,24 30 0,17 30 0,26 33 0,21 502 0,20 57 0,31 171 0,24 79 0,35 541 0,37 156 0,51 124 0,20 16 0,19
full screen enabled 2 0,06 2 0,10 3 0,01 1 0,00 3 0,05 2 0,10 2 0,07 3 0,11 3 0,07 3 0,10 2 0,06 1 0,00
Storage: quota 3650 0,60 198 0,67 12 0,03 18 0,03 23271 0,81 743 0,93 772 0,39 20 0,02 4664 0,74 91 0,21 1787 0,62 3 0,06
navigator properties 72 0,18 5 0,12 25 0,24 37 0,20 97 0,16 18 0,21 44 0,17 29 0,27 120 0,22 24 0,25 31 0,10 29 0,25
Plugins 1 0,00 1 0,00 2 0,08 2 0,08 51 0,05 4 0,01 8 0,02 21 0,17 47 0,09 4 0,09
Cookie enabled 2 0,00 2 0,01 2 0,08 2 0,07 2 0,00 2 0,01 2 0,00 2 0,07 2 0,00 2 0,07
MIME type 1 0,00 1 0,00 2 0,08 2 0,08 21 0,05 4 0,01 7 0,02 7 0,12 15 0,09 6 0,10
Time zone offset 20 0,15 7 0,12 13 0,13 19 0,15 24 0,08 9 0,17 15 0,15 14 0,11 14 0,09 9 0,14 15 0,13 12 0,18
Canvas 1360 0,46 70 0,45 90 0,31 311 0,35 619 0,33 134 0,57 247 0,39 236 0,47 1078 0,50 182 0,55 380 0,37 102 0,31
Fonts 14 0,05 2 0,06 13 0,10 17 0,11 8359 0,56 361 0,71 1934 0,52 78 0,24 1382 0,36 299 0,70 2 0,00 7 0,11
Bluetooth availability 3 0,06 2 0,05 1 0,00 1 0,00 3 0,08 3 0,10 3 0,08 1 0,00 3 0,10 1 0,00
WebGL Extensions 86 0,23 15 0,20 10 0,15 17 0,17 52 0,10 13 0,16 16 0,05 23 0,29 126 0,31 30 0,23 57 0,22 8 0,12
Audio formats: AACP 1 0,00 1 0,00 2 0,00 2 0,00 2 0,00 1 0,00 1 0,00 2 0,00 1 0,00 3 0,02
Audio formats: ACC 2 0,00 1 0,00 3 0,00 3 0,00 3 0,00 1 0,00 2 0,00 3 0,01 2 0,00 3 0,01 1 0,00 2 0,01
Audio cxt: base latency 116 0,22 9 0,22 8 0,06 11 0,05 18 0,02 3 0,03 9 0,09 13 0,11 76 0,20 5 0,04 70 0,25 7 0,05
Audio cxt: max channel count 1 0,00 1 0,00 4 0,02 4 0,03 5 0,02 5 0,02 14 0,01 7 0,03 5 0,00 5 0,02
Audio cxt: sample rate 4 0,01 2 0,00 6 0,05 7 0,04 10 0,03 5 0,04 9 0,08 10 0,09 7 0,07 3 0,09 2 0,01 6 0,04
Audio cxt: state 2 0,05 2 0,03 2 0,00 2 0,00 2 0,06 2 0,09 2 0,07 2 0,00 2 0,07 2 0,00
WebGL (Rend - Param) 525 0,46 76 0,52 10 0,09 18 0,11 2038 0,47 310 0,72 319 0,46 52 0,24 2360 0,65 137 0,54 234 0,40 7 0,05
9.2. Attribute value range ($|S|$)
@@
However, although cardinality is crucial to determine an attribute’s capacity to fingerprint individual users, we find examples where it might be partially misleading. For example, the attribute ”Screen: pixel left” can take potentially (2074, 286, 101) different values in (Chrome, Safari, Firefox) for desktop devices, respectively. In practice, (77,3%, 85,3%, 56,7%) of the samples in our fingerprint-browsers dataset take just a unique value (0) for this attribute in (Chrome, Safari, Firefox), respectively. Therefore, the discrimination power of an attribute to define unique fingerprints depends not only on its cardinality but on the frequency of appearance of its potential values.
9.3. Value frequency by attribute ($h$)
As proposed by Eckersley (eckersley2010, ), we use the entropy (masisi2008entropy, ) to measure the frequency with which different values of a given attribute appear. The higher the entropy of an attribute, the more distributed the use of different values of such an attribute is.

### v3

@@ VI Vulnerability & Accuracy
VI-A Vulnerability Metrics
VI-A1 Theoretical Vulnerability ($\mathcal{TV}$)
@@
VI-A2 Measurable Vulnerability ($\mathcal{MV}$)
@@
VI-B adF system’s accuracy
Our adF system essentially tackles a classification task, aiming to extract device fingerprints and determine their uniqueness. To assess performance, we use a standard classification metric, accuracy ($\mathcal{A}$). Accuracy, in our scenario, represents the proportion of fingerprints correctly classified as unique (true positives) or non-unique (true negatives) by our system, relative to the total number of fingerprints ($N_{f}$). A high accuracy suggests that even without access to advertising ID information, our system can effectively assess the vulnerability of different device configurations to web fingerprinting.
@@
VII Results
@@ VII-A Performance of our adF system
VII-B Vulnerability of browser-based device configurations to web fingerprinting
The values of $\mathcal{TV}$ and $\mathcal{MV}$ presented in Figure 2 serve to assess the vulnerability of each of the browser-based configurations to web fingerprinting. Given the similarity between $\mathcal{TV}$ and $\mathcal{MV}$, we discuss results based solely on $\mathcal{MV}$ for the remainder of this section.
The first conclusion we extract from the results in Figure 2 is that mobile devices exhibit lower vulnerability to web fingerprinting when compared to desktop devices for a given browser (Chrome or Safari). For instance, in desktop devices running macOS, Safari’s $\mathcal{MV}$ is 5.5%, while in mobile devices running iOS, it is 3.4%. Likewise, Chrome offers lower $\mathcal{MV}$ in mobile devices compared to desktops, regardless of the OS.
@@ VII-B Vulnerability of browser-based device configurations to web fingerprinting
Finally, when examining browsers individually, we find that Chrome ranks as the least secure option (except in the case of iOS). In contrast, Safari and Firefox provide notably superior protection against web fingerprinting. Note that we gather a few hundred samples from Firefox on both Windows and macOS, suggesting Firefox’s strong performance also on these operating systems. To better understand the outstanding performance of Chrome on iOS, we refer the reader to our analysis in Section X. This result is due to the low cardinality and entropy of attributes accessible in Chrome for iOS devices, which does not occur in other OSes.
@@
In the mobile ecosystem, 71.01% and 28.30% of the devices run on Android and iOS, respectively. Regarding browsers, 64.37% of devices employ Chrome, and 24.98% opt for Safari. By combining OSes and browsers’ market share, we identify three configurations that cover 95% of the mobile devices’ market share: {Chrome, Android} $\sim$67.7%, {Safari, iOS} $\sim$25%, and {Chrome, iOS} $\sim$3.3%. When we extrapolate the $\mathcal{MV}$ values presented in Figure 2 to this market share, we observe that web fingerprinting is feasible for at least 40% of mobile devices when exposed to ads.
In the context of the desktop ecosystem, in terms of OSes, Windows, macOS, and Linux cover 72.09%, 16.85%, and 2.77% of the market share, respectively. When it comes to browsers, Chrome dominates with 65.59%, Edge follows with 10.66%, Safari has 10.60%, and Firefox stands at 6.95%. Therefore, the most prevalent desktop configurations are: {Chrome, Windows} $\sim$54.5%, {Edge, Windows} $\sim$10.7%, {Safari, macOS} $\sim$10.6%, {Chrome, macOS} $\sim$5.7%, {Chrome, Linux} $\sim$1.2%, and {Firefox, Linux} $\sim$1.2%. When extrapolating the results from Figure 2 to this specific market share, we find that the prevalence of the {Chrome, Windows} configuration leads to a situation where at least 66% of desktop devices can be effectively fingerprinted from ads.
VII-C Mobile apps
@@
Second, Android apps are more exposed to web fingerprinting compared to iOS. In particular, when considering $\mathcal{MV}$, we find that 40.7% of Android users and 22% of iOS users can be uniquely fingerprinted using mobile apps’ ads. If we extrapolate these results to the actual market share of Android and iOS (71.01% and 28.30%, respectively), we conclude that 35% of mobile devices can be uniquely identified using web fingerprinting techniques in mobile apps.
@@
VIII Methodology Limitations
Our dataset is limited in device coverage, so our claim of fingerprint uniqueness is confined to our dataset’s scope, and we cannot claim the global uniqueness of such fingerprints. This limitation stems primarily from our restricted budget for ad campaigns, amounting to 410.27 €. Also, ethical considerations led us to exclude almost half of the samples without an advertising ID in our dataset. Nevertheless, despite these constraints, our dataset’s size (see Section VII) is comparable to previous works in the literature [gomezBoix2018hiding, andriamilanto2021largescale]. It confirms that adF provides high accuracy, establishing it as a reliable system for evaluating device configuration vulnerability to web fingerprinting. Furthermore, our results highlight the significant variability in protection among different device configurations.
We also acknowledge that our methodology faces different issues that impede collecting the required data from every single ad executing our Data Collector’s adTag. Some of these issues include: browsers running ad-blockers or disabling JavaScript, outdated browsers, limited browser API support, network problems in the connection between the ad and our server, etc. Moreover, the runtime required by our adTag represents another source of losses. Our adTag runs as long as the web-page (mobile app) where the ad is rendered is open on the user’s device. Hence, once the user closes the web-page (mobile app), the adTag execution is stopped. Callejo et al. [callejo2017opportunities] estimated that 75% of ads displayed on PCs and mobile devices remain active for at least 15 and 8 seconds, respectively. Although our adTag has an overall estimated runtime in the order of hundreds of milliseconds (depending on the browser and OS), we must also consider the rendering time of the website itself. All the mentioned causes lead to an average loss of 19% ad impressions from web-pages and of 33% ad impressions from mobile apps by comparing the campaign’s number of impressions reported by the DSP and the total samples collected in our back-end server (See Table I).
IX Causes of Web Fingerprinting Vulnerability
@@
IX-A Reported attribute count ($|N|$)
@@
Longer vectors generally result in a higher probability of acquiring unique fingerprints. In other words, device configurations that provide a greater number of attributes are typically more vulnerable to web fingerprinting. However, this is not the sole factor influencing the resilience to web fingerprinting. For instance, in the case of desktop configurations, if we use Safari’s anti-fingerprint settings in Chrome samples on Windows (i.e., we do not consider in Chrome those attributes not reported by Safari), the $\mathcal{MV}$ ($\mathcal{TV}$) for {Chrome, Windows} devices become 53.71% (60.35%). This represents an important improvement compared to the original vulnerability measured for {Chrome, Windows} configuration, 95.74% (95.81%). However, it is still far from {Safari, macOS} results ($\mathcal{MV}$ = 5.48%, $\mathcal{TV}$ = 8.36%). This example shows that two device configurations reporting exactly the same attributes present significantly different web fingerprinting vulnerabilities. Hence, there must be other factors influencing the web fingerprinting vulnerability.
Table II: Discrimination power of attributes. For each attribute and device configuration, we report: (i) whether the attribute is reported or not, (ii) the cardinality ($|S|$), (iii) the normalized entropy ($h$). We highlight in bold those attributes whose reporting we propose to block in our solution ($|S|>$ 25 and a $h\geq$ 0.10).
@@
UserAgent 6946 0.39 165 0.44 602 0.57 131 0.38 374 0.28 56 0.34 141 0.25 72 0.36 216 0.21 82 0.46 $50$ 0.73 106 0.33
CPU cores 8 0.02 2 0.01 4 0.03 4 0.05 24 0.14 9 0.22 13 0.14 4 0.12 24 0.13 9 0.23 6 0.02 2 0.03
Device memory 6 0.09 4 0.10 1 0.00 1 0.00 6 0.08 5 0.09 4 0.03 1 0.00 6 0.11 1 0.00 5 0.10 1 0.00
Screen: color depth 1 0.00 1 0.00 2 0.02 3 0.02 4 0.00 3 0.01 3 0.08 2 0.08 4 0.00 2 0.02
Screen: pixel left 4 0.00 1 0.00 1 0.00 2 0.00 1615 0.16 74 0.17 735 0.21 286 0.15 609 0.17 101 0.32
Screen: orientation angle 4 0.05 3 0.05 4 0.08 4 0.10 4 0.00 4 0.01 2 0.00 5 0.14 4 0.05 1 0.00 4 0.00 5 0.06
Screen: orientation type 4 0.05 3 0.05 4 0.08 4 0.10 4 0.00 4 0.00 3 0.00 5 0.14 4 0.07 1 0.00
Battery status: charging 3 0.04 3 0.05 1 0.00 1 0.00 3 0.03 3 0.05 3 0.08 1 0.00 3 0.07 1 0.00
Simultaneous touch points 4 0.02 1 0.00 3 0.00 4 0.00 19 0.02 6 0.05 2 0.00 3 0.08 8 0.09 2 0.00 2 0.03 2 0.01
Media devices 39 0.17 6 0.01 5 0.08 17 0.10 66 0.14 8 0.16 8 0.05 12 0.10 43 0.21 28 0.21 22 0.18 5 0.08
Languages 2748 0.31 22 0.13 27 0.15 39 0.18 1638 0.20 55 0.24 822 0.36 43 0.18 1355 0.39 40 0.25 285 0.17 42 0.20
PDF viewer enabled 3 0.01 2 0.05 3 0.08 3 0.08 3 0.04 3 0.01 3 0.01 3 0.09 3 0.08 3 0.09
User Permissions state 196 0.13 1 0.00 6 0.18 22 0.20 113 0.13 21 0.17 58 0.15 19 0.19 155 0.22 16 0.25
available height 456 0.32 41 0.31 30 0.29 32 0.25 549 0.26 79 0.42 582 0.52 310 0.45 590 0.44 206 0.63 227 0.33 20 0.24
available left 1 0.00 1 0.00 1 0.00 3 0.00 221 0.04 17 0.04 199 0.07 30 0.01 166 0.14 60 0.27
available top 1 0.00 1 0.00 1 0.00 3 0.00 484 0.03 13 0.02 181 0.18 13 0.11 170 0.12 38 0.28
available width 445 0.24 30 0.17 30 0.26 33 0.21 502 0.20 57 0.31 171 0.24 79 0.35 541 0.37 156 0.51 124 0.20 16 0.19
full screen enabled 2 0.06 2 0.10 3 0.01 1 0.00 3 0.05 2 0.10 2 0.07 3 0.11 3 0.07 3 0.10 2 0.06 1 0.00
Storage: quota 3650 0.60 198 0.67 12 0.03 18 0.03 23271 0.81 743 0.93 772 0.39 20 0.02 4664 0.74 91 0.21 1787 0.62 3 0.06
navigator properties 72 0.18 5 0.12 25 0.24 37 0.20 97 0.16 18 0.21 44 0.17 29 0.27 120 0.22 24 0.25 31 0.10 29 0.25
Plugins 1 0.00 1 0.00 2 0.08 2 0.08 51 0.05 4 0.01 8 0.02 21 0.17 47 0.09 4 0.09
Cookie enabled 2 0.00 2 0.01 2 0.08 2 0.07 2 0.00 2 0.01 2 0.00 2 0.07 2 0.00 2 0.07
MIME type 1 0.00 1 0.00 2 0.08 2 0.08 21 0.05 4 0.01 7 0.02 7 0.12 15 0.09 6 0.10
Time zone offset 20 0.15 7 0.12 13 0.13 19 0.15 24 0.08 9 0.17 15 0.15 14 0.11 14 0.09 9 0.14 15 0.13 12 0.18
Canvas 1360 0.46 70 0.45 90 0.31 311 0.35 619 0.33 134 0.57 247 0.39 236 0.47 1078 0.50 182 0.55 380 0.37 102 0.31
Fonts 14 0.05 2 0.06 13 0.10 17 0.11 8359 0.56 361 0.71 1934 0.52 78 0.24 1382 0.36 299 0.70 2 0.00 7 0.11
Bluetooth availability 3 0.06 2 0.05 1 0.00 1 0.00 3 0.08 3 0.10 3 0.08 1 0.00 3 0.10 1 0.00
WebGL Extensions 86 0.23 15 0.20 10 0.15 17 0.17 52 0.10 13 0.16 16 0.05 23 0.29 126 0.31 30 0.23 57 0.22 8 0.12
Audio formats: AACP 1 0.00 1 0.00 2 0.00 2 0.00 2 0.00 1 0.00 1 0.00 2 0.00 1 0.00 3 0.02
Audio formats: ACC 2 0.00 1 0.00 3 0.00 3 0.00 3 0.00 1 0.00 2 0.00 3 0.01 2 0.00 3 0.01 1 0.00 2 0.01
Audio cxt: base latency 116 0.22 9 0.22 8 0.06 11 0.05 18 0.02 3 0.03 9 0.09 13 0.11 76 0.20 5 0.04 70 0.25 7 0.05
Audio cxt: max channel count 1 0.00 1 0.00 4 0.02 4 0.03 5 0.02 5 0.02 14 0.01 7 0.03 5 0.00 5 0.02
Audio cxt: sample rate 4 0.01 2 0.00 6 0.05 7 0.04 10 0.03 5 0.04 9 0.08 10 0.09 7 0.07 3 0.09 2 0.01 6 0.04
Audio cxt: state 2 0.05 2 0.03 2 0.00 2 0.00 2 0.06 2 0.09 2 0.07 2 0.00 2 0.07 2 0.00
WebGL (Rend - Param) 525 0.46 76 0.52 10 0.09 18 0.11 2038 0.47 310 0.72 319 0.46 52 0.24 2360 0.65 137 0.54 234 0.40 7 0.05
IX-B Attribute value range ($|S|$)
@@
However, although cardinality is crucial to determine an attribute’s capacity to fingerprint individual users, we find examples where it might be partially misleading. For example, the attribute ”Screen: pixel left” can take potentially (2074, 286, 101) different values in (Chrome, Safari, Firefox) for desktop devices, respectively. In practice, (77.3%, 85.3%, 56.7%) of the samples in our fingerprint-browsers dataset take just a unique value (0) for this attribute in (Chrome, Safari, Firefox), respectively. Therefore, the discrimination power of an attribute to define unique fingerprints depends not only on its cardinality but also on the frequency of appearance of its potential values.
IX-C Value frequency by attribute ($h$)
As proposed by Eckersley [eckersley2010], we use the entropy to measure the frequency with which different values of a given attribute appear. The higher the entropy of an attribute, the more distributed the use of different values of such an attribute is.